Solving Exponential Equations Practice:

  1. 23x = 215
  2. 42x-1 = 4x+2
  3. 49x = 43
1
Expert's answer
2021-05-18T14:10:53-0400

"1: 2^{3x=2^{15}}"

since bases are same so power are equal

"3x=15"

"x=5"


"2: 4^{2x-1} = 4^{x+2}"

since bases are same so power are equal

"2x-1=x+2"

"x=3"


"3: 4^{9x} = 4^{3}"

since bases are same so power are equal

"9x=3"

"x=\frac{1}{3}"